OpenStudy (anonymous):

Solve. 3x^2 – 10x + 5 = 0 x=2+_ sqrt 20 x=-10 +_ sqrt 40 / 6 x=4 +_ sqrt 40 x=5 +_ sqrt 10 / 3

OpenStudy (anonymous):

x = 1/3 (5+sqrt(10)) x = 1/3 (5-sqrt(10)) a = 3 b = -10 c = 5 Use the quadratic formula

OpenStudy (anonymous):

let's cheat http://www.wolframalpha.com/input/?i=+3x^2+%E2%80%93+10x+%2B+5+%3D+0

OpenStudy (anonymous):

D

OpenStudy (anonymous):

thanks for the website. helped a lot.